# Inflationary Gravitational Wave Spectral Shapes as test for Low-Scale Leptogenesis

### v1

@@ VII Discussion & Conclusion
Leptogenesis at low and intermediate energy scales provides an attractive scenario for baryogenesis, linking the observed the matter-antimatter asymmetry to the neutrino sector and seesaw scale. In this analysis, we offer a probe of this scenario via observations of the stochastic GW background that originates during inflation (which is responsible for the creation of primordial density perturbation of the Universe) and propagates in the post-inflationary era to reach us at the present time. Particularly we showed that detecting a characteristic GW spectral shape will represent a smoking gun for a RHN or $\phi$ matter-dominated era in the early Universe. We present a minimal model realization of non-thermal leptogenesis, for which we found that a large fraction of the viable parameter space will be probed in future GW experiments. For the lightest RHN mass $M_{1}$ and washout parameter $K$ involving light-heavy neutrino Yukawa couplings $\lambda$, we find that $M_{1}$ in the range $100$ GeV $\lesssim M_{1}\lesssim 10^{14}$ GeV and $K$ for extremely weak washout ($K\sim 10^{-20}$) and for strong washout ($K\sim 10^{4}$) regime consistent with successful baryogenesis, novel GW shapes will be probes by U-DECIGO, DECIGO, BBO, $\mu-$ARES, LISA, CE, ET etc. By estimating signal-to-noise ratio (SNR) for GW experiments, we investigated the effect of the scalar mass $M_{\phi}$ and decay temperature $T_{\phi}$ which depends on the $\phi-N$ Yukawa couplings $y_{N}$.
@@
  • For vanishing Yukawa coupling $y_{R}$ defined in the Lagrangian (19), we obtain non-thermal leptogenesis which can be probed in future GW experiments such as U-DECIGO, BBO etc. (see spectra in Fig. 5, 7 and SNR in Fig. 10, 11, 12, 13).
  • Thermal leptogenesis with two-step entropy injection is possible with a non-zero $y_{R}$ (Eq. (83) and Fig. 14, 15). We propose the two-step entropy injection transfer function via Eq. (57) (see spectra in Fig. 8 and 9). Such two step will be detected in U-DECIGO, BBO, $\mu-$ARES etc. for $n_{T}=0$ and LISA, ET and CE as well for $n_{T}=0.5$.
@@
  • If $T_{N_{1}}<T_{\phi}$, then lower values of $M_{1}$ and $K$ reduce SNR and therefore challenging to test for all experiments (Fig. 10, 11, 12, 13, 14, 15).
@@
  • If $T_{\phi}<T_{N_{1}}$ in the non-thermal scenario, a lower $T_{\phi}$ value decreases the SNR and frequency of suppression $f_{\rm sup}$ given in Eq. (55), making it more difficult to observe (see spectra in Fig. 5,6). However, suitable low values of $T_{\phi}$ increases the overall GW sensitive region in the parameter space of the lightest RHN (compare Case (a) regions in Fig. 10, 11 and 12). The Case (a) scenario is discussed in Sec. IV.
@@
  • A higher $M_{\phi}$ means larger entropy injection which decreases the overall SNR values for all experiments (see spectra in Fig. 5, 6, 7, 9). $M_{\phi}$ also sets the upper bound on $M_{1}$ in our model, i.e. $M_{1}\leq M_{\phi}/2$. We are interested in the parameter space where $\phi$ is long lived in order to dominate the energy Budget if the Universe (see discussion following Eq. (25) and Eq. (29)).
@@
  • In Case (a) described in Sec. IV, where RHNs decay instantaneously after production from $\phi$ decay, the leptogenesis scale is $\sim T_{\phi}$ which means GW experiments can probe leptogenesis even for strong washout $K>1$ where RHNs do not dominate the energy budget of the Universe (Fig. 10, 11, 12, 13). This is unlike thermal leptogenesis Berbig:2023yyy where the RHN domination criterion $K\lesssim 10^{-4}$ given in Eq. (33) must be satisfied for GW sensitivity.
@@
  • Table 2 and 3 summarize which key experiments can probe thermal and non-thermal leptogenesis for $n_{T}=0$ and $0.5$ respectively, given benchmark values of the scalar mass $M_{\phi}$ and scalar decay temperature $T_{\phi}$.
@@
If the characteristic features of the GW spectral shapes proposed in this study are observed one may look to target additional observations in order to distinguish between a RHN or scalar $\phi$ dominated pre-BBN era and other forms of early matter domination. Particularly in low-scale leptogenesis RHN masses of GeV-TeV could be searched in typical Heavy Neutral Lepton Searches (HNL) search experiments (see Beacham:2019nyx ; Abdullahi:2022jlv ; Middleton:2022dio for current experimental limits) at the particle physics laboratories. In this manner we can complement GW searches with laboratory searches in the same BSM parameter space. however such a study is beyond the reach of the present paper and we plan to explore this in future. RH neutrinos, if they exist should also show up in experiments such as neutrino-less double beta decay Dolinski:2019nrj ; Gomez-Cadenas2023 or lepton number violating processes Li:2021fvw ; Cai:2017mow and therefore provide us with a myriad of pathways to independently verify the existence of an early RHN-domination or scalar domination era. This leads to a unique and exciting opportunity to form synergies between GW searches and laboratory and CMB searches.
@@ Acknowledgements.
ZAB is supported by a DST-INSPIRE fellowship. LM is supported by a UGC fellowship. Authors thank Prof. Kai Schmidt for his valuable comments on the NANOGrav best fit.
Appendix A Change of Variable
@@
$E_{\phi,N_{i}}=\rho_{\phi,N_{i}}a^{3},\quad\tilde{N}_{i}=n_{i}a^{3}=\frac{\rho_{N_{i}}}{M_{i}}a^{3},\quad\tilde{N}_{\text{B-L}}=n_{\text{B-L}}a^{3},\quad R=\rho_{R}a^{4},\quad S=s\,a^{3}$ (68)
where $a$ is the scale factor of the Universe. Instead of time $t$, we write the Boltzmann equations in terms of the ratio of scale factor $a$ to its initial value $a_{I}$, given by,
@@
$H=\sqrt{\frac{8\pi(a_{I}E_{\phi}\,y+a_{I}E_{N_{1}}\,y+a_{I}E_{N_{2}}\,y+R)}{3M_{\text{pl}}^{2}\,a_{I}^{4}\,y^{4}}}$ (70)
@@
However this is not the actual temperature of the Universe during the early matter dominated epoch. The maximum possible temperature can be approximated as,
$T_{\text{tot}}=\left[\frac{30\,(R+E_{\phi}+E_{N_{1}}+E_{N_{2}})}{\pi^{2}g_{\ast}a_{I}^{4}y^{4}}\right]^{\frac{1}{4}}$ (72)
@@
$E_{N_{i}}^{eq}$ $=\rho_{N_{i}}^{eq}a^{3}=\frac{a_{I}^{3}M_{i}^{4}y^{3}}{\pi^{2}}\left[\frac{3}{z^{2}}K_{2}(z)+\frac{1}{z}K_{1}(z)\right],\quad\tilde{N}_{i}^{eq}=\frac{E_{N_{i}}^{eq}}{M_{i}},$
$n_{N_{i}}^{eq}$ $=\frac{E_{N_{i}}^{eq}}{M_{i}y^{3}},\quad n_{R}^{eq}=\frac{3\zeta(3)}{4\pi^{2}}\times 2\left(\frac{M_{1}}{z(y)}\right)^{3}$ (74)
@@
$\Gamma_{ID}=\frac{1}{2}\sum_{i=1}^{2}\Gamma_{N_{i}}\frac{n_{N_{i}}^{eq}}{n_{R}^{eq}}$ (75)
Now we can write the Boltzmann equations given in Eq. (30) in terms of the comoving energy densities as,
@@
$\frac{dE_{N_{1}}}{dy}$ $=\frac{\Gamma_{\phi\rightarrow N_{1}N_{1}}}{\mathcal{H}}\frac{E_{\phi}}{y}-\frac{\Gamma_{N_{1}}}{\mathcal{H}y}\left(E_{N_{1}}-E_{N_{1}}^{eq}\right),$
$\frac{dE_{N_{2}}}{dy}$ $=\frac{\Gamma_{\phi\rightarrow N_{2}N_{2}}}{\mathcal{H}}\frac{E_{\phi}}{y}-\frac{\Gamma_{N_{2}}}{\mathcal{H}y}\left(E_{N_{2}}-E_{N_{2}}^{eq}\right),$ (76)
$\frac{d\tilde{N}_{B-L}}{dy}$ $=-\sum_{i=1}^{2}\frac{\Gamma_{N_{i}}}{\mathcal{H}y}\left[\epsilon\left(\tilde{N}_{i}-\tilde{N}^{eq}_{i}\right)+\frac{1}{2}\frac{n_{N_{i}}^{eq}}{n_{R}^{eq}}\tilde{N}_{B-L}\right],$
$\frac{dR}{dy}$ $=\frac{\Gamma_{\phi\rightarrow R}}{\mathcal{H}}E_{\phi}+\sum_{i=1}^{2}\frac{\Gamma_{N_{i}}}{\mathcal{H}}\left(E_{N_{i}}-E_{N_{i}}^{eq}\right)$
@@
Appendix B Thermalization of RHNs
For extremely weak washout $K\ll 1$, i.e. small light-heavy Yukawa couplings $\lambda$, RHNs can never attain thermal equilibrium with the radiation bath via decay and inverse decay. Here we show that in presence of a non-vanishing coupling $y_{R}$ with radiation, they can be thermalized through the 2-to-2 scattering process $N_{i}N_{i}\leftrightarrow\overline{f}f$ via $\phi$. We are interested in thermalization of $N_{1}$ during the $\phi-$dominated epoch. The necessary condition is,
$\Gamma_{\rm SC}\gtrsim\mathcal{H}(T)$ (78)
where $\Gamma_{\rm SC}$ is the scattering width of the 2-to-2 scattering process and $\mathcal{H}(T)\sim\frac{T^{4}}{M_{\rm pl}T_{\phi}^{2}}$ is the Hubble parameter at temperature $T$ during the $\phi-$dominated epoch Mambrini:2021zpp ; Garcia:2020eof . The scattering rate can be approximated from the scattering cross section $\sigma_{c}$ Ghoshal:2023phi ,
$\Gamma_{\rm SC}=\sigma_{\rm c}\,|v_{\rm rel}|\,n^{\rm eq}$ (79)
where $v_{\rm rel}$ is the relative velocity between the projectile and target particles and $n^{\rm eq}$ is the equilibrium number density of RHNs. Here we assumed that the cross section is velocity independent. During $\phi$ dominated era, $T\gg m_{f}$ where $m_{f}$ is the mass of the fermion. Hence fermions are relativistic with energy $\sim T$. We also assume the RHNs produced from $\phi$ are relativistic as well. Since $\phi$ is non-thermal, its tempertature need not be same as the radiation temperature, nevertheless we approximate it to be $\mathcal{O}(T)$, such that $|v_{\rm rel}|=c=1$ and $n^{\rm eq}\sim T^{3}$ Dolgov:2009zj . The total cross-section in the center-of-mass frame of the scalar is given by,
[…]
where $E_{1}\sim E_{2}\sim T$ are the energies of the incident particles and $\mathcal{M}$ is the Feynmann aplitude with over-line implying average over spin or polarization states. The integral is performed over solid angle $\Omega$. We have considered the magnitude of the 3-momenta of the incoming and outgoing particles to be equal. At tree level we get,
$\overline{|\mathcal{M}|^{2}}\sim\left(y_{R}\,y_{N_{1}}\right)^{2}$ (81)
Then the scattering rate is given by,
$\Gamma_{\rm SC}\sim\left(y_{R}\,y_{N_{1}}\right)^{2}T$ (82)
Therefore Eq. 78 becomes,
$y_{R}\,y_{N_{1}}\gtrsim\sqrt{\frac{T^{3}}{M_{\rm pl}\,T_{\phi}^{2}}}$ (83)
Taking $T=T_{\phi}$ in the above equation, one obtains the approximate condition for the RHNs to be in thermal equilibrium till $T_{\phi}$. Then at a lower temperature, the RHNs decouples and freezes out until their decay. For example, for RHNs to freeze out at around $T_{\phi}\sim 10^{5}$ GeV we need $y_{R}\,y_{N_{1}}\sim 10^{-7}$.
The two-step entropy injection scenario is possible when Eq. (83) along with the condition for $\phi$ domination given in (29) are satisfied simultaneously. Solving both of them simultaneously, we obtain the condition on scalar mass for the two-step entropy injection scenario,
$M_{\phi}\gtrsim 2\sqrt{T_{\phi}\,M_{\rm pl}}+M_{\rm pl}\left(y_{N_{1}}-y_{R}\right)^{2}\gtrsim 2\sqrt{T_{\phi}\,M_{\rm pl}}$ (84)
where the last inequality is the minimum requirement for $M_{\phi}$ for which thermal scenario is possible when $y_{N_{1}}\sim y_{R}$.

### v2

@@ VII Discussion & Conclusion
Leptogenesis at low and intermediate energy scales provides an attractive scenario for baryogenesis, linking the observed matter-antimatter asymmetry to the neutrino sector and seesaw scale. In this analysis, we offer a probe of this scenario via observations of the stochastic GW background that originates during inflation (which is responsible for the creation of primordial density perturbation of the Universe) and propagates in the post-inflationary era to reach us at present. Particularly, we constructed a global $U(1)_{B-L}$ model of neutrino mass that facilitates non-thermal leptogenesis through the decay of the Goldstone boson called majoron, denoted by $\phi$. We showed that detecting a characteristic GW spectral shape will provide an indication of a RHN or $\phi$ matter-dominated era in the early Universe. We present a minimal model realization of non-thermal leptogenesis, for which we found that a large fraction of the viable parameter space will be probed in future GW experiments. For the lightest RHN mass $M_{1}$ and washout parameter $K$ involving light-heavy neutrino Yukawa couplings $\lambda$, we find that $M_{1}$ in the range $100$ GeV $\lesssim M_{1}\lesssim 10^{14}$ GeV and $K$ for extremely weak washout ($K\sim 10^{-20}$) and for strong washout ($K\sim 10^{4}$) regime consistent with successful baryogenesis, novel GW shapes will be probes by U-DECIGO, DECIGO, BBO, $\mu$-ARES, LISA, CE, ET etc. By estimating signal-to-noise ratio (SNR) for GW experiments, we investigated the effect of the majoron mass $M_{\phi}$ and decay temperature $T_{\phi}$ which depends on the $\phi-N$ Yukawa couplings $y_{N}$.
@@
  • Non-thermal leptogenesis can be probed in future GW experiments such as U-DECIGO, BBO etc. (see spectra in Fig. 8-9 and SNR in Fig. 10 - 13).
@@
  • If $T_{N_{1}}\ll T_{\phi}$ in Case (c) scenario with 2-step entropy injection, lower values of $M_{1}$ and $K$ reduce SNR and therefore challenging to test for all experiments (Fig. 10-13).
@@
  • For Case (a) and Case (b) scenarios discussed in Sec. IV, a lower $T_{\phi}$ value decreases the SNR and frequency of suppression $f_{\rm sup}$ given in Eq. (68), making it more difficult to observe (see spectra in Fig. 8). However, suitable low values of $T_{\phi}$ increase the overall GW sensitive region in the parameter space of the lightest RHN (compare for example Fig. 10 and 11).
@@
  • A higher $M_{\phi}$ in general means larger entropy injection which decreases the overall SNR values for all experiments (compare solid and dashed spectra in Fig. 8). However, for Case (c), higher $M_{\phi}$ increases the SNR (solid and dashed spectra in Fig. 9 (Right)). $M_{\phi}$ also sets the upper bound on $M_{1}$ in our model, i.e. $M_{1}\leq M_{\phi}/2$. We are interested in the parameter space where $\phi$ is long-lived to dominate the energy Budget of the Universe (see discussion following Eq. (34) and Eq. (42)).
@@
  • In Case (a) and Case (b) described in Sec. IV, the leptogenesis scale is $\sim T_{\phi}$ which means GW experiments can probe leptogenesis even for strong washout $K>1$ where RHNs might eventually thermalize with the radiation bath (Fig. 10-13). This is unlike the model in Berbig:2023yyy where the RHN domination criterion $K\lesssim 4\times 10^{-4}$ must be satisfied for GW sensitivity.
@@
  • Table 2 summarizes which key experiments can probe non-thermal leptogenesis for $n_{T}=0$ and $0.3$ respectively, given benchmark values of the majoron mass $M_{\phi}$ and decay temperature $T_{\phi}$.
@@
If the characteristic features of the GW spectral shapes proposed in this study are observed, one may look to target additional observations to distinguish between an RHN or majoron $\phi$ dominated pre-BBN era and other forms of early matter domination. Particularly in low-scale leptogenesis RHN masses of GeV-TeV could be searched in typical Heavy Neutral Lepton Searches (HNL) search experiments (see Beacham:2019nyx ; Abdullahi:2022jlv ; Middleton:2022dio for current experimental limits) at the particle physics laboratories. In this manner, we can complement GW searches with laboratory searches in the same BSM parameter space. However, such a study is beyond the reach of the present paper and we plan to explore this in future. RH neutrinos, if they exist should also show up in experiments such as neutrino-less double beta decay Dolinski:2019nrj ; Gomez-Cadenas2023 or lepton number violating processes Li:2021fvw ; Cai:2017mow and therefore provide us with a myriad of pathways to independently verify the existence of an early RHN-domination or majoron domination era. This leads to a unique and exciting opportunity to form synergies between GW searches and laboratory and CMB searches.
@@ Acknowledgements.
ZAB is supported by a DST-INSPIRE fellowship. LM is supported by a UGC fellowship. The authors thank Prof. Kai Schmitz for his valuable comments.
Appendix A Structure of complex orthogonal matrix $R$
Here we explicitly show that CP asymmetry defined in Eq. (18) can be large even for a tiny washout parameter $K$ defined in Eq. (10). For example, take a simple form of $R$ defined in Eq. (11),
$R=\begin{pmatrix}\cos\omega&0&-\sin\omega\\ 0&1&0\\ \sin\omega&0&\cos\omega\\ \end{pmatrix}$, (81)
where $\omega=x+iy$ is a complex angle. Assuming normal ordering and $m_{1}=0$, the CP asymmetry can be written in terms of $x$ and $y$ as,
$|\epsilon_{1}|=\frac{3}{32\pi}\frac{M_{1}m_{3}}{v^{2}}\frac{\sin 2x\sinh 2y}{\sin^{2}x\cosh^{2}y+\cos^{2}x\sinh^{2}y}=\frac{3}{16\pi}\frac{M_{1}m_{3}}{v^{2}}\frac{\sin 2x\sinh 2y}{\cosh 2y-\cos 2x}.$ (82)
When $x=y\ll 1$, it can be verified that $|\epsilon_{1}|=\epsilon_{1}^{\rm max}$, which is the maximal CP asymmetry defined in Eq. (19). Taking $x=y=10^{-11}$ gives $|\epsilon_{1}|\sim 10^{-6}$ (assuming $M_{1}=10^{10}$ GeV) while $K\sim 10^{-20}$. In Fig. 14 we show the regions where $|\epsilon_{1}|>10^{-7}$ (blue) and $|\epsilon_{1}|>7\times 10^{-7}$ (orange) in the parameter space spanned by $x$ and $y$. Along with it, we also show the contours of constant $K$ values in the same parameter space. For this plot we assume normal ordering and $m_{1}=0$, $M_{1}=10^{10}$ GeV, $M_{2}=10^{11}$ GeV, $M_{3}=10^{12}$ GeV. We see that for such high RHN mass scales, it is possible to generate $|\epsilon_{1}|=\epsilon_{1}^{\rm max}$ even if $\tilde{m}_{1}\rightarrow 0$ hence $K\rightarrow 0$. However $K_{2},K_{3}$ remains large.
Figure 14: CP parameter $|\epsilon_{1}|$ defined in Eq. (18) and contours of decay parameter $K$ defined in Eq. (10) in the paramter space spanned by $x$ and $y$. Blue and orange region signifies $|\epsilon_{1}|>10^{-7}$ and $7\times 10^{-7}$ respectively. It is seen that for $x=y\ll 1$, large $|\epsilon_{1}|$ can be obtained even for a tiny value of $K$.
Appendix B Change of Variable
@@
$E_{\phi,N_{i}}=\rho_{\phi,N_{i}}a^{3},\quad\tilde{N}_{1}=n_{1}a^{3}=\frac{\rho_{N_{1}}}{M_{1}}a^{3},\quad\tilde{N}_{\text{B-L}}=n_{\text{B-L}}a^{3},\quad R=\rho_{R}a^{4},\quad S=s\,a^{3}$ (83)
where $a$ is the scale factor of the Universe. Note that $E_{N_{1}},\tilde{N}_{1}$ become comoving quantities only after RHNs become non-relativistic, i.e. at temperatures below $T_{\rm NR}$ defined in Eq. (37). Instead of time $t$, we write the Boltzmann equations in terms of the ratio of scale factor $a$ to its initial value $a_{I}$, given by,
@@
$H=\sqrt{\frac{8\pi(a_{I}E_{\phi}\,y+a_{I}E_{N_{1}}\,y+R)}{3M_{\text{pl}}^{2}\,a_{I}^{4}\,y^{4}}}$ (85)
@@
$E_{N_{1}}^{eq}$ $=\rho_{N_{1}}^{eq}a^{3}=\frac{a_{I}^{3}M_{1}^{4}y^{3}}{\pi^{2}}\left[\frac{3}{z^{2}}K_{2}(z)+\frac{1}{z}K_{1}(z)\right],\quad\tilde{N}_{1}^{eq}=\frac{E_{N_{1}}^{eq}}{M_{1}},$
$n_{N_{1}}^{eq}$ $=\frac{E_{N_{1}}^{eq}}{M_{1}y^{3}},\quad n_{R}^{eq}=\frac{3\zeta(3)}{4\pi^{2}}\times 2\left(\frac{M_{1}}{z(y)}\right)^{3}$ (88)
@@
$\Gamma_{ID}=\frac{1}{2}\Gamma_{N_{1}}\frac{n_{N_{1}}^{eq}}{n_{R}^{eq}}$ (89)
As described earlier, we ignore the inverse decay term of $\phi$ as well as the RHN scattering terms involving $\Gamma_{\rm scatt}$ and $\Gamma_{ID}^{\rm scatt}$ in the Boltzmann equations given in Eq. (43), and rewrite them as
@@
$\frac{dE_{N_{1}}}{dy}$ $=\frac{\Gamma_{\phi\rightarrow N_{1}N_{1}}}{\mathcal{H}}\frac{E_{\phi}}{y}-\frac{\Gamma_{N_{1}}}{\mathcal{H}y}\left(E_{N_{1}}-E_{N_{1}}^{eq}\right)-\frac{E_{N_{1}}}{y}\Theta(T-T_{\rm NR}),$
$\frac{d\tilde{N}_{B-L}}{dy}$ $=-\frac{\Gamma_{N_{1}}}{\mathcal{H}y}\left[\epsilon\left(\tilde{N}_{1}-\tilde{N}^{eq}_{1}\right)+\frac{1}{2}\frac{n_{N_{1}}^{eq}}{n_{R}^{eq}}\tilde{N}_{B-L}\right],$
$\frac{dR}{dy}$ $=\frac{\Gamma_{\phi\rightarrow R}}{\mathcal{H}}E_{\phi}+\frac{\Gamma_{N_{1}}}{\mathcal{H}}\left(E_{N_{1}}-E_{N_{1}}^{eq}\right)$
where $\Theta(x)$ is the Heaviside step function separating the relativistic stage of RHNs from the non-relativistic stage,
$\Theta(x)=\begin{cases}1&\text{for }x>1\\ 0&\text{otherwise}\end{cases}$ (90)